McBride and Associates employs two professional appraisers, each having a different specialty. Debbie specializes in commercial appraisals and Tara specializes in residential appraisals. The company expects to incur total overhead costs of $492,030 during the year and applies overhead based on annual salary costs. The salaries and billable hours of the two appraisers are estimated to be as follows:

Debbie Tara
Annual Salary $ 222,720 $ 105,300
Billable Hours 2,900 2,250

The accountant for McBride and Associates is computing the hourly rate that should be used to charge clients for Debbie and Tara’s services. The hourly billing rate should be set to cover the total cost of services (salary plus overhead) plus a 20 percent markup.

Required:
1.
Compute the predetermined overhead rate.



2. Compute the hourly billing rate for Debbie and Tara. (Do not round your intermediate calculations, and round your final answers to 2 decimal places.)
Answer #1

1. Calculation of Predetermined overhead rate

  Total overhead cost = Debbie's Salary+ Tara's Salary = $(222720+105300)= $328020

Expected Overhead Cost= $492030

Predetermined Overhead Rate= $( 492030/328020)= $1.5

2. Calculation of Hourly Billing Rate:

  Debbie's Salary= $222720

Overhead cost= $(222720*1.5)= $334080

Total = $(222720+334080) =$556800

Per Hour Billing cost= $(556800/2900)= $192 per hr.

20% Mark up cost= 192*20% = $38.4 per hr

Billing Rate= $(192+38.4)= $230.4 per hr or $230 per hr

Tara's Salary= $105300

Overhead cost= $(105300*1.5)= $157950

Total = $(105300+157950) =$263250

Per Hour Billing cost= $(263250/2250)= $117 per hr.

20% Mark up cost= 117*20% = $23.4 per hr

Billing Rate= $(117+23.4)= $140.4 per hr or $140 per hr